A=B-C,B=C-D,C=D-A AND ABCD IS NOT EQUAL TO 0.WHAT IS VALUE OF(A/B)+(B/C)+(C/D)+(D/A)

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
A=B-C ---------(1)
       B=C-D ---------(2)
       C=D-A ---------(3)
Adding (1) , (2) and (3)
     C= - 2A
Putting this value of C in (1)
We get B= -A
And outing the value of C in (3)
We get D= -A
 
Now put these values of B, C and D in terms of A in the required expression to get
 
A/-A + A/2A + 2A/A + -A/A
= ½
 
Answer = 1/2 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)